### v1

# GrAviPaSt’s Lens to the Past: Unveiling the Evolution of Filamentary Structures

Abstract
This paper examines the evolution of cosmic filaments across redshifts 1, 0.5, and 0 using the IllustrisTNG100-1 magneto-hydrodynamical simulation. To achieve this, we introduce GrAviPaSt, a simple and efficient filament identification method that leverages gravitational potential, an A*-like path-finding algorithm, and spanning trees. Applying this method to galaxy distributions at different redshifts allows us to analyze various filament properties, including their length, thickness, mass density contrast, and radial profile. Additionally, we investigate dynamic characteristics such as the mean distance of filament galaxies from the skeleton, their weighted mean velocity, and velocity trends normalized by their positions within the filaments. Our findings reveal the evolution of cosmic filaments from redshift 1 to 0, highlighting key differences across classifications. In particular, we examine the mass density contrast radial profile of filaments connecting two galaxy groups and those linking two galaxy clusters, identifying distinct differences in profile shape between these categories. Furthermore, in the context of weighted mean velocity, we analyze cosmic filaments exhibiting either negative or positive weighted mean velocity, demonstrating their differing evolutionary trends in terms of the mean distance of filament galaxies from the skeleton.
[…]
show]P.Ghafour@outlook.com ]tavasoli@ipm.ir
1 Introduction
The large-scale structure (LSS) of the universe reveals an intricate network known as the cosmic web, composed of clusters (or knots), filaments, walls (or sheets), and vast low-density regions known as cosmic voids (M. Davis et al. (1982); J. R. Bond et al. (1996); P. J. E. Peebles (2020)). Observations from redshift surveys (e.g., V. De Lapparent et al. (1986); M. Colless et al. (2001)) and insights from numerical simulations (e.g., V. Springel et al. (2005); M. Vogelsberger et al. (2014)) contributed significantly to identifying the LSS.
The cosmic web originates from primordial density fluctuations that grew through gravitational instability, causing dark matter to collapse under gravitational forces and forming the filamentary structures observed in the galaxy distribution today (P. Peebles (1967); Y. B. Zel’dovich (1970); Y. B. Zeldovich et al. (1982)). Within this structure, matter flows from less dense environments like voids to denser regions such as walls, filaments, and nodes, where it accumulates (D. Galárraga-Espinosa et al. (2022)).
Galaxies within the cosmic web are not randomly distributed; their properties, such as mass, activity, morphology and luminosity, are closely correlated with their environment (e.g., T. Kuutma et al. (2017); P. Wang et al. (2018)). Characterizing the topology, density and dynamical state of the cosmic web is therefore crucial for understanding the evolution of galaxies and refining cosmological models (I. Santiago-Bautista et al. (2020)).
Cosmic filaments are fundamental structures of the LSS, forming a crucial part of the cosmic web and acting as boundaries for voids while bridging galaxy groups and clusters across intergalactic space (M. Cautun et al. (2014); D. Zakharova et al. (2023); W. Wang et al. (2024)). These elongated features span vast distances, channeling matter along their lengths toward galaxy halos and clusters and playing a pivotal role in shaping the universe’s LSS (Y. B. Zeldovich et al. (1982); E. Tempel et al. (2014)).
The filamentary environment significantly influences the evolution of galaxies. Factors such as local environmental density and proximity to the filament’s central axis play critical roles in influencing galaxy development (V. Bonjean et al. (2020); P. Ganeshaiah Veena et al. (2021)). Studies have demonstrated that galaxies closer to the filament skeleton exhibit higher fractions of passive evolution compared to those situated farther away in the surrounding void regions, highlighting the pivotal role of filaments in regulating galaxy properties and star formation activity (e.g., A. Singh et al. (2020)).
Filaments not only host galaxies composed of cooler baryonic matter but also serve as conduits for warm-hot intergalactic matter, significantly contributing to the cosmic baryon cycle (D. Galárraga-Espinosa et al. (2022)). Also, some large-scale phenomena within filaments, such as cosmic web enhancement and detachment, have been identified and linked to the presence of ionized gas clouds in the circumgalactic environments of certain filament galaxies (B. Vulcani et al. (2019); M. A. Aragon-Calvo et al. (2019)). Furthermore, extensive research has investigated the statistical characteristics of filaments, including their lengths, radial density profiles and magnetic field strengths, providing deeper insights into their structure and impact (N. Malavasi et al. (2020); T. Vernstrom et al. (2021); T. Yang et al. (2022); D. Galárraga-Espinosa et al. (2022)). Studies on hydrodynamical simulations have further illustrated how filaments act as conduits connecting galaxies to clusters, refining our understanding of the complex processes underlying cosmic evolution (e.g., A. Rost et al. (2021); C. Gouin et al. (2022)).
Observations of filaments have primarily been derived from galaxy redshift surveys, such as the Sloan Digital Sky Survey (SDSS; K. N. Abazajian et al. (2009)) and the Dark Energy Spectroscopic Instrument (DESI; A. Adame et al. (2024)). These observations reveal individual filaments spanning several megaparsecs, forming bridges of matter connecting pairs of galaxy clusters (e.g., V. Biffi et al. (2022); A. D. Hincks et al. (2022)). Meanwhile, cosmological simulations provide a more detailed understanding of their intricate geometry and their function in channeling matter from voids to galaxy clusters (e.g., Z. Wang et al. (2024)). Despite their significant role in galaxy formation and cosmic evolution, the complex and elongated shapes of filaments present challenges for their identification in both observational data and simulations based on dark matter particles (D. Zakharova et al. (2023)).
Over the years, a wide range of methods have been developed to identify cosmic filaments by analyzing the spatial distribution of particles in simulations and observational galaxy data. These approaches include geometric and topological evaluations, such as the Bisous model (E. Tempel et al. (2016)), Skeleton (T. Sousbie et al. (2008)), NEXUS (M. Cautun et al. (2013)) and tessellation-based tools like DisPerSE (T. Sousbie et al. (2011)). Other techniques, including multiscale morphology filters MMF (M. A. Aragón-Calvo et al. (2007, 2010)), segmentation-based models like the Candy model (R. S. Stoica et al. (2005)), and watershed methods such as SpineWeb (M. A. Aragón-Calvo et al. (2010)), offer diverse ways to detect filaments. Each of these methods has unique strengths, yet comparing their effectiveness is complicated by the absence of a universally accepted definition of cosmic filaments and the varying requirements across studies (D. Alina et al. (2022); S. Inoue et al. (2022); Y. Zhang et al. (2022); M. Aragon-Calvo (2024)).
The results produced by filament detection algorithms are highly sensitive to the choice of parameters and identification methods. For example, smoothing scales significantly influence the width and length distributions of filaments identified in density maps, as observed in methods applied to surveys such as SDSS and DEEP2 (N. A. Bond et al. (2010); E. Choi et al. (2010)). Similarly, algorithms like DisPerSE demonstrate how parameters like persistence thresholds can alter filament properties, removing shorter filaments or modifying length distributions (N. Malavasi et al. (2020)). Additionally, the process of determining filament endpoints plays a crucial role, as this directly impacts length measurements (N. Malavasi et al. (2020)). Statistical analyses also reveal that filament characteristics can differ based on the types of tracers used, whether galaxies or dark matter particles.
An effective filament detection algorithm should first and foremost align with human visual perception while providing quantitative and reproducible results rooted in sound numerical theory. Furthermore, it should minimize dependency on adjustable parameters, ensuring reliable outcomes with simplicity. Efficiency is also a critical factor, as the algorithm must achieve high performance without excessive computational demands (I. Santiago-Bautista et al. (2020)).
Finally, studies have highlighted intriguing differences in the radial density profiles of filaments. For example, investigations using the DisPerSE algorithm on simulations like Illustris-TNG show that shorter filaments exhibit distinct density profiles compared to their longer counterparts (D. Galárraga-Espinosa et al. (2022)).
In this study, we present a novel yet simple and effective method for detecting cosmic filaments. This method integrates Gravitational potential, an A*-like Path-finding algorithm, and Spanning trees (GrAviPaSt) to identify filamentary structures within galaxy distributions as described in Section 3. We then apply the GrAviPaSt method to galaxy distributions at three distinct redshifts: 0, 0.5, and 1, and present the results in Section 4. By comparing these findings across different redshifts, we investigate the evolutionary trends of these cosmic filaments over cosmic time. The galaxy distributions used in this study are sourced from the IllustrisTNG100-1 simulation, as detailed in Section 2. Finally, in Section 5, we summarize our findings and further discuss the results presented in Section 4.
2 Data: Illustris-TNG
The dataset for this study is derived from the gravo-magnetohydrodynamical IllustrisTNG [^1] https://www.tng-project.org simulation (D. Nelson et al. (2019)), which simulates the coupled evolution of dark matter, gas, stars, and black holes throughout cosmic history, spanning redshift z = 127 to z = 0. This simulation employs the Arepo moving-mesh framework (V. Springel (2010)) and adopts cosmological parameters informed by the Planck cosmology (P. A. Ade et al. (2016)). These include density parameters of $\Omega_{m,0}$ = 0.3089 for matter, $\Omega_{b,0}$ = 0.0486 for baryons, and $\Omega_{\Lambda,0}$ = 0.6911 for the cosmological constant, along with $n_{s}$ = 0.9667 for the scalar spectral index, $\sigma_{8}$ = 0.8159 for the amplitude of fluctuations, and $H_{0}=100h$ $(km/s/Mpc)$ with $h_{0}$ = 0.6774 for the Hubble constant.
For this analysis, the focus is on the IllustrisTNG100-1 simulation box, which represents the highest-resolution run within the IllustrisTNG100 suite. This box encompasses a cubic volume with a side length of approximately 75 $(Mpc/h)$, featuring a dark matter and baryonic resolution of $7.5\times 10^{6}M_{\odot}$ and $1.4\times 10^{6}M_{\odot}$ respectively. This combination of spatial scale and resolution allows for both the reliable identification of extended cosmic filaments and a detailed analysis of their diverse characteristics.
In this study, we selected all simulated galaxies from the illustrisTNG100-1 catalog that are brighter than approximately -16 in the r-band filter, possess stellar masses exceeding $10^{8}$ $M_{\odot}$, and exist at three different redshifts. We used the snapshots 50, 67 and 99, respectively corresponding to redshifts 1.0, 0.5 and 0. Using these thresholds, our final samples in these three redshifts contains about $34000$, $42000$ and $42000$ galaxies respectively.
3 Methodology: GrAviPaSt
In this section, we introduce a novel, simple and yet efficient cosmic filament-finding algorithm called the GrAviPaSt method. This algorithm is specifically designed to operate with a minimal number of parameters and ensuring computational efficiency. Due to its simplicity, the GrAviPaSt method can be applied to large datasets while requiring a reasonable amount of computational resources.
The GrAviPaSt method combines several powerful techniques to achieve its goals. It leverages the minimum spanning tree (MST) algorithm (D. R. Karger et al. (1995)), integrates the concept of gravitational potential, utilizes an A*-like path-finding algorithm (P. E. Hart et al. (1968)), and takes into account the inherent geometry of the input dataset. This method consists of four main phases, which are applied sequentially to the input dataset. Each phase involves a series of steps.
3.1 Prim’s MST
In the first phase, GrAviPaSt method employs the minimum spanning tree (MST) algorithm to identify pairs of galaxy groups and search for the micro-filament structures (hereafter micro-filaments represent filaments connecting two groups of galaxies) between them.
GrAviPaSt method specifically uses Prim’s MST algorithm (R. C. Prim (1957)) instead of other MST algorithms such as Kruskal’s algorithm (e.g., as mentioned in I. Santiago-Bautista et al. (2020)). This choice is due to the optimal implementation of Prim’s MST. In GrAviPaSt method, the spatial coordinates of each galaxy group’s BCG (or BGG) serve as the node coordinates for the input of the MST algorithm. The Prim’s MST algorithm begins by initializing the process with an arbitrary node (vertex) from the input dataset, designating it as part of the growing minimum spanning tree (MST), while assigning an initial cost of infinity to all other nodes except the starting node, which has a cost of zero. Next, it iteratively selects the edge with the smallest weight that connects any node in the MST to a node outside it. For each newly added node, the algorithm updates the costs of its neighboring nodes, ensuring that any smaller edge weight replaces the current cost. This process is repeated until all nodes are included in the MST, thereby creating a structure with minimum-cost connections. Following the implementation of Prim’s MST, the pairs of galaxy groups between which the GrAviPaSt method will search for micro-filaments are identified. An example of the implementation of Prim’s MST is shown in Figure 1.
Figure 1: A schematic illustration depicting the sequential phases and steps of the GrAviPaSt method in identifying a macro-filament composed of three micro-filaments. Four galaxy groups are presented, with the method’s phases applied sequentially from bottom to top.
3.2 A*-like Path-Finder
In the second phase, the GrAviPaSt method employs an A*-like path-finding algorithm combined with gravitational potential. This phase consists of multiple sequential steps applied to the input dataset. To implement this phase, the GrAviPaSt method utilizes the output of the MST algorithm from the previous phase.
For each pair of galaxy groups determined by the MST algorithm (Section 3.1), the GrAviPaSt method first identifies the minimum and maximum coordinates of the galaxies within these two groups. It then crops the dataset box to these coordinates, creating a rectangular cuboid (hereafter referred to as the cropped box), which contains only the galaxies inside this defined space. It is worth noting that the GrAviPaSt method includes all galaxies within the cropped box, not just those belonging to the two galaxy groups.
Next, the GrAviPaSt method centers two spheres on the coordinates of the BCGs (or BGGs) of the two groups, using the virial radius $R_{200}$ of each group as the radius of the spheres. Subsequently, it tangentially places two pairs of planes around these spheres. Each pair of planes is parallel to each other but perpendicular to the other pair of planes. Using these four planes, the GrAviPaSt method constructs a pyramid-shaped structure around the two spheres, with the line connecting the two BCGs (or BGGs) forming the central axis. The algorithm retains only the region that is inside the pyramid, outside the spheres of the galaxy groups, and between the two spheres. The resulting space is a truncated pyramid: its base plane is tangential to the sphere of the larger group, its top plane is tangential to the sphere of the smaller group, and its four trapezoidal sides are defined by the two pairs of planes.
With the truncated pyramid and cropped box in place, the GrAviPaSt method moves to the next step of this phase. In this step, the truncated pyramid is filled with grid points, each separated by a distance of $d$ Mpc from its horizontal and vertical neighbors. The value of d is determined by:
$d=h\times(\frac{N}{V})^{\frac{1}{3}}$ (1)
where $h$ ($H_{0}/100$) is the Hubble constant, $N$ is the total number of galaxies in the input dataset (not the cropped box), and $V$ is the total volume of the input dataset. These grid points are shown between the galaxy group pairs in Figure 1.
Once the grid points are established, the GrAviPaSt method proceeds to implement the A*-like path-finding algorithm. This involves several steps, including determining the starting point, selecting the ending point candidates, and identifying the path with the minimum gravitational potential to locate the deepest gravitational well between two groups.
To determine the starting point, the GrAviPaSt method identifies grid points adjacent to the top plane of the truncated pyramid as candidates. Thus, the starting point is always selected near the smaller galaxy group. For each starting point candidate, the gravitational potential of the galaxies within the cropped box is calculated using:
$\Phi=-G\sum_{i}\frac{M_{i}}{\sqrt{(x-x_{i})^{2}+(y-y_{i})^{2}+(z-z_{i})^{2}}}$ (2)
where the coordinates $x$, $y$, and $z$ correspond to the starting point candidate, and the parameters with the $i$-subscript refer to all galaxies within the cropped box. Among these candidates, the grid point with the lowest (deepest) gravitational potential is chosen as the starting point.
For the ending point candidates, the GrAviPaSt method considers grid points adjacent to the base plane of the truncated pyramid, ensuring that the end of the path is always located near the larger galaxy group. With both the starting point and the ending point candidates established, the A*-like path-finding algorithm can now be applied.
To identify the path with the deepest gravitational potential between two groups, the A*-like path-finding algorithm begins at the starting point. From this point, it considers all neighboring grid points that are closer to the BCG (or BGG) of the larger group than to the starting point. For each of these neighboring grid points, GrAviPaSt calculates the gravitational potential using Equation 2 and selects the grid point with the lowest (deepest) gravitational potential as the next step in the path. This process is repeated iteratively, with each subsequent point added to the path, until the algorithm reaches one of the ending point candidates. Once an ending point candidate is reached, the path-finding process halts.
After determining this minimum path (or deepest gravitational well), the GrAviPaSt method considers this path as the main skeleton of the micro-filament connecting the two galaxy groups. These skeletons are illustrated between galaxy groups in Figure 1.
3.3 Micro-Filaments Structure
In the third phase, GrAviPaSt method constructs the final structure of the micro-filaments based on the distribution of galaxies surrounding the skeleton (deepest potential path) identified between the two groups of galaxies in the previous phase.
The GrAviPaSt method models micro-filaments as cylindrical structures that bridge the two galaxy groups, with their central axis aligning with the skeleton identified in the previous phase (Section 3.2). To determine the radius of these micro-filaments, GrAviPaSt examines the galaxies surrounding the skeleton. Specifically, it considers galaxies that lie outside the virial radius of the groups, between the two groups, and within a distance from the skeleton that is less than or equal to the average of the two groups’ virial radius $R_{200}$. Among these galaxies, the GrAviPaSt method identifies the galaxy farthest from the skeleton and assigns this distance as the radius of the micro-filament.
Once the final structure of the micro-filaments is determined, the GrAviPaSt method evaluates all the micro-filaments and eliminates those containing only a single galaxy. As a result, all of the final micro-filaments, as well as the macro-filaments (Section 3.4), contain at least two galaxies within their cylindrical structure. This cylindrical structure is depicted in Figure 1.
3.4 Macro-Filaments
The final phase involves grouping the micro-filaments like chains to construct the macro-filaments. To achieve this, the GrAviPaSt method follows a three-step process based on the thickness, angle, and gravitational potential of the micro-filaments associated with each node, determining which pairs of micro-filaments connected to a node can be grouped into the same macro-filament.
First, the algorithm calculates the $2\sigma$ (standard deviation) of all micro-filament thicknesses connected to a node, identifying the thickness error for that node. The GrAviPaSt method begins with the thickest micro-filament connected to the node as the base micro-filament and considers other micro-filaments with thicknesses within the $\pm$ error range of the base micro-filament as pair candidates.
Second, among these pair candidates, the algorithm calculates the angle between each candidate and the base micro-filament. To do so, the GrAviPaSt method uses the coordinates of the starting or ending points of the pair candidates corresponding to the node. By connecting these points to the node, the algorithm calculates the angle between the candidates and the base micro-filament. It always considers the smaller angle between the pair candidates and the base micro-filament. The GrAviPaSt method then retains only those pair candidates with an angle greater than 90 degrees, removing the others from consideration.
In the final step, if multiple pair candidates meet the thickness and angle criteria, the GrAviPaSt method uses gravitational potential to determine the best pair. It calculates the difference between the gravitational potential of the points of the pair candidates connected to the node and that of the base micro-filament. The algorithm selects the micro-filament with the lowest potential difference as the best pair. If no suitable pair is found for the base micro-filament, the algorithm moves to next thickest micro-filament connected to the node and considers it as the base micro-filament, repeating the process until a pair is identified or all micro-filaments associated with the node are examined without finding a pair. After determining the best pair for each node, the algorithm groups paired micro-filaments together like a chain, forming the final structure of the macro-filaments. These macro-filaments can consist of one or more micro-filaments.
4 Results
This section presents the findings derived from the application of the GrAviPaSt method to the three datasets described in Section 2. By identifying both micro- and macro-filaments using this method, we analyze their various characteristics across different redshifts: 1, 0.5, and 0, covering approximately the last 8 Gyrs.
For the Prim’s MST phase of the GrAviPaSt method, the coordinates of BGGs (or BCGs) are chosen from galaxy groups containing at least 10 members. These selected coordinates serve as the nodes for the MST algorithm, with the micro-filaments identified in this study forming bridges between them.
4.1 Micro-Filaments
In this subsection, we examine the characteristics of the micro-filaments identified using the GrAviPaSt method. First, we analyze the length and thickness distributions of these filaments, providing insights into their structural properties. Subsequently, we investigate the respected velocity of filament galaxies along the micro-filament skeleton, revealing their dynamic behavior. Finally, we explore the radial profiles of the micro-filaments to identify patterns in their density and structural composition.
4.1.1 Length Evolution
One of the primary characteristics analyzed is the length distribution of micro-filaments across different redshifts. The length of these micro-filaments is determined by summing the distances between neighboring grid points that form the micro-filament skeleton (Section 3.2).
The distribution of micro-filament lengths is illustrated in Figure 2, with the top panel showing the co-moving lengths, and middle panel the proper lengths. As shown in this figure, the co-moving lengths of micro-filaments remain within a similar range across different redshifts, exhibiting slow evolution. By factoring out the Hubble flow, the specific influences of gravity and DE on filament length evolution become more apparent (D. Galárraga-Espinosa et al. (2024)).
In contrast, the distribution of proper filament lengths varies significantly across different redshifts. As illustrated in the middle panel of Figure 2, the range of micro-filament lengths expands with decreasing redshift. This global growth, evident in proper coordinates, is a direct consequence of the Universe’s expansion. Driven by the Hubble flow, the cosmic web undergoes continuous stretching, leading to an overall increase in micro-filament lengths.
Figure 2: The median trend of micro-filament thickness, along with its associated error, is presented for their length distribution in co-moving (Top) and proper (Middle) coordinates. At the Bottom, the median trend of the velocity of galaxies along the micro-filaments’ skeleton is shown. Note that, according to the implementation of the GrAviPaSt method, the skeleton of micro-filaments originates at the smaller group and terminates at the larger group.
The statistical analysis, performed using two-sided Kolmogorov–Smirnov (K-S) tests, further supports these findings. Comparisons of the co-moving length distributions of micro-filaments at redshifts 0 and 0.5, 0 and 1, as well as 0.5 and 1, yield p-values of 0.92, 0.68, and 0.52, respectively. These values indicate no statistically significant differences among the distributions. In contrast, comparisons of the proper length distributions result in p-values of $1.8\times 10^{-5}$, $9.34\times 10^{-15}$, and $1.21\times 10^{-4}$ respectively, indicating statistically significant differences among these distributions. Additionally, Figure 2 illustrates that the micro-filament population, in both proper and co-moving coordinates, is consistently dominated by shorter filaments across all redshifts.
4.1.2 Thickness Evolution
The thickness distribution of micro-filaments across different redshifts, in both co-moving and proper coordinates, represents another key characteristic examined in this study. As illustrated in the panels of Figure 2, the range of thickness distributions in co-moving coordinates remains largely consistent across redshifts, similar to the length distributions. However, in proper coordinates, as redshift decreases, the range of thickness distributions expands. This thickening of micro-filaments over cosmic time, akin to the trend observed in proper length distributions, is driven by the Hubble flow and the Universe’s continuous expansion.
Further statistical analysis using K-S tests reinforces these findings. Comparisons of co-moving thickness distributions at redshifts 0 and 0.5, 0 and 1, and 0.5 and 1 yield p-values of 0.99, 0.23, and 0.22, respectively, indicating no statistically significant differences among the distributions, consistent with the observations for micro-filament length distributions. Conversely, comparisons in proper coordinates result in p-values of $8.15\times 10^{-10}$, $1.88\times 10^{-39}$, and $3.57\times 10^{-13}$, demonstrating statistically differences among these distributions.
4.1.3 Filament Galaxies Respected Velocity
As another characteristic of micro-filaments, we examine the velocities of filament galaxies relative to the micro-filament skeleton, accounting for their normalized positions along the micro-filament. To achieve this, we identify the nearest grid point on the micro-filament skeleton for each galaxy. The length of the micro-filament from its starting point to this grid point is calculated (Section 4.1.1) and then normalized by the total length of the micro-filament to determine the galaxy’s normalized position.
It is important to note that the micro-filament skeletons, as identified by the GrAviPaSt method, begin at the smaller galaxy group and extend to the larger one (Section 3.2). Smaller normalized positions correspond to galaxies near the smaller group, while larger normalized positions indicate galaxies closer to the larger group. In the bottom panel of Figure 2, negative velocities represent galaxies moving toward the micro-filament skeleton, while positive velocities indicate galaxies moving away from it. As shown in this figure, galaxies with smaller normalized positions, and therefore closer to the smaller group, tend to exhibit more negative velocities and approaching the filament skeleton. Moving further along the micro-filament skeleton, galaxy velocities gradually become more positive. Within the normalized range of 0 to 0.3, the median velocities are negative across different redshifts, transitioning to values close to zero around 0.3. Beyond this point, within the range of 0.3 to 1, the median velocities remain approximately zero.
4.1.4 Radial Profiles
The final attribute examined for micro-filaments is their radial profile. Given that the GrAviPaSt method identifies micro-filaments based on the gravitational potential of galaxies (Section 3), we explored the variations in mass density contrast along their radii. To derive this profile, we first measured the perpendicular distance of each filament galaxy from the micro-filament skeleton and then normalized it to the micro-filament’s radius. With these normalized distances, we then combined all filament galaxies, stacking them to create a comprehensive micro-filament that integrates galaxies across different radial positions. The mass density contrast was then determined at various normalized radial distances from the skeleton by:
$\delta_{m}(r)=\frac{\rho_{m}(r)-\rho_{m}^{b}}{\rho_{m}^{b}}$ (3)
where $\rho_{m}(r)$ represents the mass density at a normalized radius $r$, $\delta_{m}(r)$ denotes the corresponding mass density contrast, and $\rho_{m}^{b}$ refers to the background mass density.
To model and compare these radial profiles across different redshifts, we applied the generic-$\beta$ model from the Plummer profile (A. Cavaliere & R. Fusco-Femiano (1976); M. Arnaud (2009); S. Ettori et al. (2013); D. Galárraga-Espinosa et al. (2022)). This profile is expressed as:
$1+\delta_{m}(r)=\frac{1+\delta_{m}^{1}}{(1+r^{\alpha})^{\beta}}$ (4)
where $\delta_{m}(r)$ is from Equation 3, and $\delta_{m}^{1}$ represents the mass density contrast at $r=1$. Given that our analysis focuses on logarithmic scales, we reformulated this profile as:
$log_{10}(\delta_{m}(r))=log_{10}\left(\frac{1+\delta_{m}^{1}}{(1+r^{\alpha})^{\beta}}-1\right)+\gamma$ (5)
where $\alpha$, $\beta$, and $\gamma$ serve as fitting parameters.
We selected this profile due to its effective representation of the total mass density contrast of micro-filaments. The radial mass density contrast profile of micro-filaments across three different redshifts is illustrated in the top panel of Figure 3, with the corresponding fit-curves shown as dashed lines. The fitted parameters and Mean Squared Errors (MSEs) are summarized in Table 1. As depicted in this figure, the mass density contrast declines as the distance from the micro-filament skeleton increases. Additionally, as redshift decreases, micro-filaments exhibit a gradual reduction in mass density contrast, reflecting a slow evolutionary trend.
Figure 3: The radial profiles of mass density contrast are presented for total (top), CC (middle) and GG micro-filaments (bottom), each displayed across three redshifts using different colors. The corresponding fit curves, based on the Plummer profile, are represented by dashed lines.
Beyond the total population of micro-filaments, we also examined two specific types: those connecting galaxy clusters (CCs) and those linking groups of galaxies (GGs). CC micro-filaments were identified as structures bridging galaxy groups with at least 100 members, while GG micro-filaments connected groups containing 10 to 20 members. The middle and bottom panels of Figure 3 illustrate the profiles and corresponding fit-curves for CC and GG micro-filaments, modeled using the Plummer profile. Their fitted parameters and MSEs are detailed in Table 1.
As shown in Figure 3, the radial mass density profiles of CC micro-filaments across all three redshifts exhibit a general shape similar to that of the overall micro-filament population. In contrast, the profiles of GG micro-filaments differ significantly from those of both total filaments and CC micro-filaments.
Table 1: The fitted parameters of the Plummer profile, as defined in Equation 5, along with the Mean Squared Errors (MSEs) of the corresponding fit curves.
Redshift z=0 z=0.5 z=1
Parameter $\alpha$ $\beta$ $\gamma$ $MSE$ $\alpha$ $\beta$ $\gamma$ $MSE$ $\alpha$ $\beta$ $\gamma$ $MSE$
Total 0.79 4.81 1.41 0.0005 0.82 4.49 1.32 0.0002 0.56 6.97 2.07 0.0005
CCs 0.87 5.56 1.70 0.0015 0.47 6.91 2.38 0.0007 0.77 7.27 2.27 0.0044
GGs 0.37 5.79 1.69 0.0098 0.14 6.89 2.25 0.0019 0.16 7.38 2.49 0.0053
4.2 Macro-Filaments
This subsection examines the properties of macro-filaments identified by the GrAviPaSt method. As outlined in Section 3.4, this method links micro-filaments using a three-step procedure that considers their thickness, angular alignment, and gravitational potential. In this analysis, we explore the characteristics of these macro-filaments and their evolution across different redshifts. Specifically, we investigate their length distribution, mass density contrast, and the average distance between filament galaxies and the skeleton of the macro-filaments. These parameters are expressed in proper coordinates rather than co-moving coordinates to incorporate the effects of Hubble flow and the expansion of the Universe.
4.2.1 Effective Length Evolution
The effective length (hereafter referred to as length) of macro-filaments is calculated as the cumulative sum of the lengths of all micro-filaments that constitute the macro-filament, while not considering the diameters of the galaxy groups within it.
As observed, the range of length distributions differs across redshifts, with macro-filaments gradually increasing in length as redshift decreases. In all three redshifts, shorter macro-filaments make up the majority of the population; however, their dominance diminishes as redshift decreases. Kernel Density Estimation (KDE) curves are overlaid on the length histograms, showing peaks at 6.86, 4.59, and 3.98 for redshifts 0, 0.5, and 1, respectively. These results indicate evolution in the length distributions over cosmic time. Additionally, a two-sided K-S test conducted on the length distributions of macro-filaments between redshifts 0 and 0.5, 0 and 1, 0.5 and 1, yields p-values of 0.06, $4.85\times 10^{-6}$, and $1.6\times 10^{-3}$ respectively, indicating a statistically significant difference between the length distributions at redshifts 0 and 0.5 compared to that at redshift 1.
Figure 4: The distribution of macro-filaments length in proper coordinates and mass density contrast across three redshifts is shown. Scatter plots and the mass density contrast histograms use distinct colors to represent the two macro-filament populations, categorized by the number of filament galaxies they host. Horizontal lines mark the median length within each population’s distribution, while vertical lines represent the median mass density contrast at each redshift.
The population of macro-filaments can be categorized in multiple ways. Based on length, the population is dominated by shorter macro-filaments, similar to the micro-filaments (see Section 4.1.1). Another categorization considers the number of filament galaxies hosted by macro-filaments, dividing them into two groups accordingly. These groups are represented by different colors in Figure 4. Macro-filaments with a low galaxy count ($\leq 10$) undergo slow evolutionary changes over cosmic time and consistently maintain shorter lengths across all redshifts. As depicted in Figure 4, both groups exhibit a gradual increase in proper length, though the gap between their median values widens as redshift decreases.
4.2.2 Mass Density Contrast Evolution
Another key characteristic of macro-filaments is their mass density contrast. This metric is preferred over the traditional number density (e.g., investigated in Y. Zhang et al. (2024); D. Galárraga-Espinosa et al. (2024)) due to the implementation of the GrAviPaSt method, which relies on the gravitational potential of the galaxy distribution to identify the structures of both micro- and macro-filaments. To investigate this property, the mass density of each macro-filament is first computed by summing the total stellar mass of the galaxies hosted within the macro-filament and dividing it by the total volume of the macro-filament. The mass density contrast is then determined using the density contrast formula (Equation 3).
As illustrated in Figure 4, the mass density contrast distribution remains nearly consistent across different redshifts for both groups. The median of each group exhibits a gradual decline as redshift decreases, primarily due to the thickening of micro-filaments, as discussed earlier (Section 4.1.2). Additionally, the gap between the medians of the two groups widens with decreasing redshift. A statistical analysis using a two-sided K-S test on the mass density contrast distributions of macro-filaments with a lower galaxy population ($\leq 10$) produced p-values of 0.74, 0.02, and 0.25 for comparisons between redshifts 0 and 0.5, 0 and 1, and 0.5 and 1, respectively. These results suggest no significant difference between the distributions at redshifts 0 and 0.5, as well as between 0.5 and 1; however, a statistically significant difference is observed between redshifts 0 and 1. For macro-filaments with a higher galaxy population ($>10$), the same K-S tests yielded p-values of 0.98, 0.79, and 0.76 for the same redshift comparisons. These findings indicate no significant differences in mass density contrast distributions of this group across the three redshifts.
Table 2: Total luminosity ($log_{10}$ $L_{\odot}$), mass ($log_{10}$ $M_{\odot}$) and volume of the macro-filaments in both co-moving ($(Mpc/h)^{3}$) and proper ($Mpc^{3}$) through different redshifts.
Redshift z=0 z=0.5 z=1
Total mass $1.74\times 10^{4}$ $1.83\times 10^{4}$ $1.54\times 10^{4}$
Total luminosity $1.75\times 10^{4}$ $1.85\times 10^{4}$ $1.56\times 10^{4}$
Total proper volume 1415.82 603.53 242.05
Total co-moving volume 440.09 428.21 423.11
4.2.3 Filament Galaxies Mean Radial Distance
The next characteristic of macro-filaments examined is the distribution of the mean distance and weighted mean velocity of filament galaxies relative to the skeleton identified by the GrAviPaSt method (Section 3.2). To compute these parameters, the perpendicular distance between each filament galaxy and the associated skeleton is measured and defined as the galaxy’s radial distance. The velocity of the galaxies is calculated following the details outlined in Section 4.1.3, with the modification of using the stellar mass of galaxies as weights for the weighted mean. For each macro-filament, the mean values of these parameters are then determined. These results are illustrated in Figure 5, plotted against the lengths of the macro-filaments.
Figure 5: The distribution of the mean distance to the skeleton of macro-filaments across different redshifts is shown with respect to the lengths of the macro-filaments. Macro-filaments with negative and positive weighted mean velocities of their hosted galaxies are depicted using distinct colors.
As shown in Figure 5, the mean distance distribution of filament galaxies expands as redshift decreases. At redshift 1, the distribution is more concentrated, but as redshift decreases, it becomes more dispersed, covering a broader range. To further analyze these dynamics, macro-filaments are classified based on the weighted mean velocity of their hosted galaxies. Two categories are defined: macro-filaments with a negative weighted mean velocity, indicating a predominance of matter infall toward the filament skeleton, and macro-filaments with a positive weighted mean velocity, signifying a predominance of outflow motion away from the skeleton. As depicted in Figure 5, the median values of both groups increase with decreasing redshift, and the gap between their medians becomes more distinct over cosmic time. The fraction of macro-filaments exhibiting a negative weighted mean velocity is 0.52, 0.51, and 0.57 for redshifts 0, 0.5, and 1, respectively. The stability of this fraction across redshifts suggests slow dynamical evolution of macro-filaments over cosmic time. Additionally, the mean of weighted mean velocities for the negative group is -153.61 $km/s$, -157.54 $km/s$, and -141.76 $km/s$ at redshifts 0, 0.5, and 1, respectively, whereas for the positive group, it is 175.64 $km/s$, 186.69 $km/s$, and 201.51 $km/s$ across these redshifts. The consistency of mean values within the negative group across different redshifts, along with the gradual decline in mean values for the positive group as redshift decreases, indicates the absorption of matter toward the filament skeletons identified by GrAviPaSt.
A two-sided K-S test was performed to analyze the distribution of mean distance to the skeleton for the negative group. The resulting p-values are $1.73\times 10^{-9}$, $9.37\times 10^{-23}$, and $7.21\times 10^{-12}$ for comparisons between redshifts 0 and 0.5, 0 and 1, and 0.5 and 1, respectively. Similarly, the K-S test applied to the positive group produced p-values of $1.33\times 10^{-4}$, $2.27\times 10^{-19}$, and $4.34\times 10^{-10}$ for the same redshift comparisons. These results demonstrate a statistically significant difference in the distribution of both groups across different redshifts.
4.2.4 Total Luminosity, Mass and Volume Evolution
Finally we investigate the evolution of the total luminosity and mass of filament galaxies hosted by macro-filaments, as well as the total volume of the macro-filaments in both the proper and co-moving coordinates. The specific details of these parameters are presented in Table 2.
From the data in Table 2, it is clear that the total volume of macro-filaments steadily increases from redshift 1 to 0 in both proper and co-moving coordinates. This growth corresponds to the thickening of micro-filaments across these redshifts (see Section 4.1.2). As expected, when accounting for the Hubble flow and the universe’s expansion in the proper volume of macro-filaments, the increase is more pronounced compared to the co-moving coordinate. Furthermore, the total mass and luminosity of filament galaxies demonstrate an overall upward trend, with a significant increase observed between redshifts 1 and 0.5. However, this is followed by a minor decline between redshifts 0.5 and 0.
5 Summary and Discussion
The analyses presented in this study revealed that both micro- and macro-filaments exhibit consistent evolution across redshifts 1, 0.5, and 0. The observed distributions of filaments length (Section 4.1.1 and Section 4.2.1), thickness (Section 4.1.2), mass density contrast (Section 4.2.2) and radial profile (Section 4.1.4) are in general agreement with previous studies (e.g., D. Galárraga-Espinosa et al. (2024, 2020)) that employed alternative filament identification algorithms, such as COWS (S. Pfeifer et al. (2022)), NEXUS (M. Cautun et al. (2013)), and DisPerSE (T. Sousbie et al. (2011)). These findings reinforce the reliability of GrAviPaSt method in identifying and characterizing cosmic filaments.
In addition, the various characteristics of the filaments provide valuable insights into their dynamics. For instance, the velocity patterns of filament galaxies along the length of micro-filaments. Moreover, the mean weighted velocity of these galaxies highlights the overall inflow and outflow of matter relative to the filament skeleton. The observations based on these weighted mean velocities indicates the absorption of matter toward the cosmic filament skeletons (Section 4.2.3). Additionally, the relatively close mean distance of galaxies to the skeleton for both categories of macro-filaments across different redshifts emphasizes the structural consistency of these filamentary structures (Section 4.2.3).
These analyses have been presented to illustrate various aspects of the cosmic filaments and their dynamics. However, further investigation on some of these characteristics is necessary in future studies. The key objectives of this analysis were to assess the effectiveness of the GrAviPaSt method in identifying filaments and to uncover new insights into the evolutionary trends of cosmic filaments from redshift 1 to 0. Nevertheless, it remains challenging to fully assess the success of any filament identification approach due to the lack of a detailed consensus on the definition of cosmic filaments, while different studies have employed diverse methods and criteria to define these structures.
In conclusion, the findings demonstrate that the evolution and characteristics of micro- and macro-filaments identified using the GrAviPaSt method are in general consistent with existing literature while also highlighting areas for future exploration to enhance the study of cosmic structures.